OpenStudy (anonymous):

A 187 foot tall building casts a shadow 38 feet long, how long is the shadow of a 6 foot tall man? Round to the nearest tenth of a foot. Answers 0.8 ft 29.5 ft 1.2 ft 0.3 ft

OpenStudy (anonymous):

(38/187) = (x/6) cross multiply 187x = 228 x=1.2 the shadow is 1.2 feet long

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem :)